A homeowner jogged from 1 of his subdivision to gate 2 in a straight 235 m track in 4.25 minutes turned around and jogged 125 m toward gate 1 in another 1.50 minute. What are the average speeda and average velocities of the home owner for (A) the first 4.25, (B) the last 1.50 min, and (C) the total time he jogged
1
Expert's answer
2020-10-15T10:50:16-0400

a)


"v=\frac{235}{4.25(60)}=0.923\frac{m}{s}""\vec{v}=+0.923\frac{m}{s}"


b)



"v=\frac{125}{1.5(60)}=1.389\frac{m}{s}""\vec{v}=-1.389\frac{m}{s}"


c)


"T=4.25(60)+1.5(60)=345\ S"